Jamil always throws loose change into a pencil holder on his desk and takes it out every two weeks. This time it is all nickels and dimes. There are twice as many dimes as nickels, and the value of the dimes is $ 0.90 more than the value of the nickels. How many nickels does Jamil have?
Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52778)

.
Let x be the number of nickles; then the number of dimes is  2x.


The "value" equation is


    10*(2x) - 5x = 90   cents.


Simplify and solve it


    20x - 5x = 90

    15x = 90

      x = 90/15 = 6.


Answer.   6 nickels  and  2*6 = 12 dimes.

Solved.


Answer by stanbon(75887)

Jamil always throws loose change into a pencil holder on his desk and takes it out every two weeks. This time it is all nickels and dimes. There are twice as many dimes as nickels, and the value of the dimes is $ 0.90 more than the value of the nickels. How many nickels does Jamil have?
Equations:
d = 2n
10d - 5n = 90 cents
------------------------------
Substitute for "d" and solve for "n"::
10(2n) - 5n = 90
15n = 90
n = 6 (# of nickels)
------------------
Cheers,
Stan H.
----------------------